if sec thita + tan thita=x find the value of sec thita

Grade:10

2 Answers

Ashmita Bhattacharyya
12 years ago

secthita+tanthita=x.......1

sec2thita-tan2thita=1

(secthita-tanthita)(secthita+tanthita)=1

(secthita-tanthita)x=1

secthita-tanthita=1/x......2

adding 1 and 2,

secthita+tanthita+secthita-tanthita=1+1/x

secthita=(x+1)/2x(Ans)

Parshant Kaushik
12 years ago

sec thita =x2+1/2x